# An analysis of Wigner’s friend in the framework of quantum mechanics based on the principle of typicality

### v1

@@
Abstract. The notion of probability plays a crucial role in quantum mechanics. It appears in quantum mechanics as the Born rule. In modern mathematics which describes quantum mechanics, however, probability theory means nothing other than measure theory, and therefore any operational characterization of the notion of probability is still missing in quantum mechanics. In our former works [K. Tadaki, arXiv:1804.10174], based on the toolkit of algorithmic randomness, we presented a refinement of the Born rule, called the principle of typicality, for specifying the property of results of measurements in an operational way. The Wigner’s friend paradox is a Gedankenexperiment regarding when and where the reduction of the state vector occurs in a chain of the measurements by several observers where the state of the consciousness of each observer is measured by the subsequent observer, except for the last observer in the chain. It is one of the central open questions in the measurement problem of quantum mechanics. In this paper, we make an analysis of the Wigner’s friend paradox in the framework of quantum mechanics based on the principle of typicality. We then draw common sense conclusions about the Wigner’s friend paradox. In addition, we make a prediction, which is testable in principle, about its variant proposed by Deutsch.
@@ 1.1 Contributions of the paper
In this paper, we extend the framework of the principle of typicality, which was introduced by Tadaki [42, 43, 45, 46, 47], so that it can be applicable to the situation, such as the Wigner’s friend paradox, where apparatuses perform measurements over other apparatuses as well as over a normal quantum system only being measured. For that purpose, we introduce postulates, Postulates 6 and 7 below, for determining the states of apparatuses in early stages among all apparatuses constitute a measurement process. We make an analysis of the Wigner’s friend paradox within this extended framework of quantum mechanics based on the principle of typicality. We then draw common sense conclusions about the Wigner’s friend paradox.
Deutsch’s thought experiment [11] is a variant of the Wigner’s friend paradox, which can, in principle, verify the effect of the consciousness of observer on the reduction of the state vector. In this paper, we make an analysis of the Deutsch’s thought experiment comprehensively within the extended framework of the principle of typicality. Based on this, we make a prediction which is testable in principle.
Based on the extended framework of the principle of typicality, we can analyze still more complicated situations than the Wigner’s friend paradox or Deutsch’s thought experiment. As such a example, we introduce the combination of the Wigner’s friend paradox and Deutsch’s thought experiment, called the Wigner-Deutsch collaboration. We make an analysis of it comprehensively in this paper.
@@ 2.1 Basic notation and definitions
Let $\mathcal{H}$ be a complex Hilbert space. A projector $P$ in $\mathcal{H}$ is a non-zero Hermitian bounded operator in $\mathcal{H}$ such that $P^{2}=P$. A collection $\{P_{a}\}_{a\in\Theta}$ of projectors in $\mathcal{H}$ is called a projection-valued measure (PVM, for short) if $\Theta$ is an alphabet and the following hold:
@@
    $\sum_{a\in\Theta}P_{a}=I$, where $I$ denotes the identity operator in $\mathcal{H}$.
@@ 2.2 Martin-Löf randomness with respect to an arbitrary probability measure
We introduce the notion of Martin-Löf randomness with respect to an arbitrary probability measure. For that purpose, we briefly review measure theory according to Nies [31, Section 1.9]. See also Billingsley [4] for measure theory in general. We here summurize the description of Tadaki [44, Section 2.2], which is based on Nies [31, Section 1.9].
@@ 4.1 Marginal distribution
Let $\Omega$ and $\Theta$ be alphabets. Let $P\in\mathbb{P}(\Omega\times\Theta)$, and let $\alpha\in(\Omega\times\Theta)^{\infty}$. Suppose that $\beta$ is an infinite sequence over $\Omega$ obtained from $\alpha$ by replacing each element $(m,l)$ in $\alpha$ by $m$. If $\alpha$ is Martin-Löf $P$-random then $\beta$ is Martin-Löf $Q$-random, where $Q\in\mathbb{P}(\Omega)$ such that
$Q(m):=\sum_{l\in\Theta}P(m,l)$
for every $m\in\Omega$. ∎
@@ 5 Postulates of quantum mechanics
In this section, we review the central postulates of (the conventional) quantum mechanics. For simplicity, in this paper we consider the postulates of quantum mechanics for a finite-dimensional quantum system, i.e., a quantum system whose state space is a finite-dimensional Hilbert space. Nielsen and Chuang [30] treat thoroughly the postulates of (the conventional) quantum mechanics in the finite-dimensional case, as a textbook of the field of quantum computation and quantum information in which such a case is typical. In this paper we refer to the postulates of the conventional quantum mechanics in the form presented in Nielsen and Chuang [30, Section 2.2].
@@
Associated to any isolated physical system is a complex vector space with inner product (that is, a Hilbert space) known as the state space of the system. The system is completely described by its state vector, which is a unit vector in the system’s state space. ∎
@@
The evolution of a closed quantum system is described by a unitary transformation. That is, the state $|\Psi_{1}\rangle$ of the system at time $t_{1}$ is related to the state $|\Psi_{2}\rangle$ of the system at time $t_{2}$ by a unitary operator $U$, which depends only on the times $t_{1}$ and $t_{2}$, in such a way that $|\Psi_{2}\rangle=U|\Psi_{1}\rangle$. ∎
@@
A quantum measurement is described by a collection $\{M_{m}\}_{m\in\Omega}$ of measurement operators which satisfy the completeness equation,
$\sum_{m\in\Omega}M_{m}^{\dagger}M_{m}=I.$
These are operators acting on the state space of the system being measured.
@@
    The set of possible outcomes of the measurement equals the finite set $\Omega$. If the state of the quantum system is $|\Psi\rangle$ immediately before the measurement then the probability that result $m$ occurs is given by
@@
In this paper as well as our former works [36, 39, 40, 42, 43, 45, 46, 47, 49, 50], we keep Postulates 1, 2, and 3 in their original forms without any modifications. The principle of typicality, Postulate 5 below, is proposed as a refinement of Postulate 4 for replacing it, based on the notion of Martin-Löf randomness with respect to a probability measure.
@@ 6 The principle of typicality: Our world is typical
Tadaki [42] reformulated the original framework of MWI by Everett [22] in a form of mathematical rigor from a modern point of view. The point of this rigorous treatment of the framework of MWI is the use of the notion of probability measure representation and its induction of probability measure, as presented in Section 2.2. Tadaki [42] then introduced the principle of typicality in this rigorous framework. The principle of typicality is an operational refinement of Postulate 4 based on the notion of Martin-Löf randomness with respect to a probability measure given in Definition 2.1. In what follows, we review this rigorous framework of MWI based the principle of typicality. See Tadaki [47] for the details of this framework of MWI and its validity.
Now, according to Tadaki [47], let us introduce the setting of MWI in terms of our terminology in a form of mathematical rigor. Let $\mathcal{S}$ be an arbitrary quantum system with state space $\mathcal{H}$ of finite dimension. Consider a measurement over $\mathcal{S}$ described by arbitrary measurement operators $\{M_{m}\}_{m\in\Omega}$ satisfying the completeness equation,
@@
Here, $\Omega$ is the set of all possible outcomes of the measurement. [^1] The set $\Omega$ is finite, and therefore it is an alphabet. Let $\mathcal{A}$ be an apparatus performing the measurement described by $\{M_{m}\}_{m\in\Omega}$, which is a quantum system with state space $\overline{\mathcal{H}}$. According to Postulates 1, 2, and 3, the measurement process of the measurement described by the measurement operators $\{M_{m}\}_{m\in\Omega}$ is described by a unitary operator $U$ such that
@@
for every $|\Psi\rangle\in\mathcal{H}$. This is due to von Neumann [53, Section VI.3] in the case where the measurement operators $\{M_{m}\}_{m\in\Omega}$ form a PVM such that $\operatorname{rank}M_{m}=1$ for every $m\in\Omega$. For the derivation of (5) in the general case, see Tadaki [47, Section 7.1]. Actually, $U$ describes the interaction between the system $\mathcal{S}$ and the apparatus $\mathcal{A}$. The vector $|\Phi_{\mathrm{init}}\rangle\in\overline{\mathcal{H}}$ is the initial state of the apparatus $\mathcal{A}$, and $|\Phi[m]\rangle\in\overline{\mathcal{H}}$ is a final state of the apparatus $\mathcal{A}$ for each $m\in\Omega$, with $\langle\Phi[m]|\Phi[m^{\prime}]\rangle=\delta_{m,m^{\prime}}$. For every $m\in\Omega$, the state $|\Phi[m]\rangle$ indicates that the apparatus $\mathcal{A}$ records the value $m$ as the measurement outcome. By the unitary interaction (5) as a measurement process, a correlation (i.e., entanglement) is generated between the system $\mathcal{S}$ and the apparatus $\mathcal{A}$.
@@
Then, for aiming at deriving Postulate 4, MWI assigns “weight” to each of worlds. Namely, it introduces a probability measure on the set of all worlds in the following manner: First, MWI introduces a probability measure representation on the set of prefixes of worlds, i.e., the set $\Omega^{*}$ in this case. This probability measure representation is given by a function $r\colon\Omega^{*}\to[0,1]$ with
@@
which is the square of the norm of each state $(M_{m_{1}}|\Psi_{1}\rangle)\otimes\dots\otimes(M_{m_{n}}|\Psi_{n}\rangle)\otimes|\Phi[m_{1}\dots m_{n}]\rangle$ in the superposition (6). Using the completeness equation (4), it is easy to check that $r$ is certainly a probability measure representation over $\Omega$. We call the probability measure representation $r$ the probability measure representation for the prefixes of worlds. Then MWI tries to derive Postulate 4 by adopting the probability measure induced by the probability measure representation $r$ for the prefixes of worlds as the probability measure on the set of all worlds.
@@
Consider an arbitrary finite-dimensional quantum system $\mathcal{S}$ and a measurement over $\mathcal{S}$ described by arbitrary measurement operators $\{M_{m}\}_{m\in\Omega}$, where the measurement process is described by (5) as an interaction of the system $\mathcal{S}$ with an apparatus $\mathcal{A}$. We suppose the following situation:
@@
    For each $n\in\mathbb{N}^{+}$, the system $\mathcal{S}_{n}$ is prepared in a state $|\Psi_{n}\rangle$, [^2] In Definition 6.1, all $|\Psi_{n}\rangle$ are not required to be an identical state. while the apparatus $\mathcal{A}_{n}$ is prepared in a state $|\Phi_{\mathrm{init}}\rangle$, and then the measurement described by $\{M_{m}\}_{m\in\Omega}$ is performed over $\mathcal{S}_{n}$ by interacting it with the apparatus $\mathcal{A}_{n}$ according to the unitary time-evolution (5).
@@
As mentioned above, the original MWI by Everett [22] aimed to derive Postulate 4 (i), the Born rule, from the remaining postulates. However, it would seem impossible to do this for several reasons (see Tadaki [47] for these reasons). Instead, it is appropriate to introduce an additional postulate in the rigorous framework of MWI developed above, in order to overcome the defect of the original MWI and to make quantum mechanics operationally perfect. Thus, we put forward the principle of typicality as follows.
@@
We assume the notation and notions introduced and defined in Definition 6.1, and consider the same setting of measurements as considered in Definition 6.1. Let $\omega\in\Omega^{\infty}$ be our world. Then $\omega$ is Martin-Löf random with respect to the probability measure $\mu_{p}$ induced by the probability measure representation $p\colon\Omega^{*}\to[0,1]$ for the prefixes of worlds. Moreover, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the composite system $\mathcal{S}_{n}+\mathcal{A}_{n}$, which consists of the $n$th copy $\mathcal{S}_{n}$ of the system $\mathcal{S}$ and the $n$th copy $\mathcal{A}_{n}$ of the apparatus $\mathcal{A}$, immediately after the measurement by $\mathcal{A}_{n}$ performed over $\mathcal{S}_{n}$ is given by
@@
7 Apparatuses measuring apparatuses and their confirming points

### v2

@@
Abstract. The notion of probability plays a crucial role in quantum mechanics. It appears in quantum mechanics as the Born rule. In modern mathematics which describes quantum mechanics, however, probability theory means nothing other than measure theory, and therefore any operational characterization of the notion of probability is still missing in quantum mechanics. In our former works [K. Tadaki, arXiv:1804.10174], based on the toolkit of algorithmic randomness, we presented a refinement of the Born rule, called the principle of typicality, for specifying the property of results of measurements in an operational way. The Wigner’s friend paradox is a Gedankenexperiment regarding when and where the reduction of the state vector occurs in a chain of the measurements by several observers where the state of the consciousness of each observer is measured by the subsequent observer, except for the last observer in the chain. It is one of the central open questions in the measurement problem of quantum mechanics. In this paper, we make an analysis of the Wigner’s friend paradox in the framework of quantum mechanics based on the principle of typicality. We then draw common sense conclusions about the Wigner’s friend paradox. Deutsch’s thought experiment is a variant of the Wigner’s friend paradox, which can, in principle, verify the effect of the consciousness of observer on the reduction of the state vector. We make an analysis of Deutsch’s thought experiment comprehensively in our framework. We then make a prediction which is testable in principle. In our framework, we can analyze still more complicated situations than the Wigner’s friend paradox or Deutsch’s thought experiment. In this paper, as such an example, we introduce a combination of these two, called the Wigner-Deutsch collaboration. We make an analysis of it comprehensively in our framework.
@@ 1.1 Contributions of the paper
In this paper, we extend the framework of the principle of typicality, which was introduced by Tadaki [42, 43, 45, 46, 47], so that it can be applicable to the situation, such as the Wigner’s friend paradox, where apparatuses perform measurements over other apparatuses as well as over a normal quantum system only being measured. For that purpose, we introduce postulates, Postulates 6 and 7 below, for determining the states of the apparatuses in early stages among all apparatuses which constitute a measurement process. We make an analysis of the Wigner’s friend paradox within this extended framework of quantum mechanics based on the principle of typicality. We then draw common sense conclusions about the Wigner’s friend paradox.
Deutsch’s thought experiment [11] is a variant of the Wigner’s friend paradox, which can, in principle, verify the effect of the consciousness of observer on the reduction of the state vector. In this paper, we make an analysis of Deutsch’s thought experiment comprehensively within the extended framework of the principle of typicality. Based on this, we make a prediction which is testable in principle.
Based on the extended framework of the principle of typicality, we can analyze still more complicated situations than the Wigner’s friend paradox or Deutsch’s thought experiment. As such an example, we introduce a combination of the Wigner’s friend paradox and Deutsch’s thought experiment, called the Wigner-Deutsch collaboration. We make an analysis of it comprehensively in this paper. This analysis shows how effectively our extended framework works in an even more complicated situation.
@@ 2.1 Basic notation and definitions
Let $\mathcal{H}$ be an arbitrary complex Hilbert space. A projector $P$ on $\mathcal{H}$ is a Hermitian bounded operator on $\mathcal{H}$ such that $P^{2}=P$. A collection $\{P_{a}\}_{a\in\Theta}$ of projectors on $\mathcal{H}$ is called a projection-valued measure (PVM, for short) in $\mathcal{H}$ if $\Theta$ is an alphabet and the following (i) and (ii) hold:
@@
    $\sum_{a\in\Theta}P_{a}=I$, where $I$ denotes the identity operator on $\mathcal{H}$.
@@ 2.2 Martin-Löf randomness with respect to an arbitrary probability measure
We introduce the notion of Martin-Löf randomness with respect to an arbitrary probability measure. For that purpose, we briefly review measure theory according to Nies [31, Section 1.9]. See also Billingsley [4] for measure theory in general. We here summarize the description of Tadaki [44, Section 2.2], which is based on Nies [31, Section 1.9].
@@ 4.1 Marginal distribution
Let $\Omega$ and $\Theta$ be alphabets. Let $P\in\mathbb{P}(\Omega\times\Theta)$, and let $\alpha\in(\Omega\times\Theta)^{\infty}$. Suppose that $\beta$ is an infinite sequence over $\Omega$ obtained from $\alpha$ by replacing each element $(l,m)$ in $\alpha$ by $l$. If $\alpha$ is Martin-Löf $P$-random then $\beta$ is Martin-Löf $Q$-random, where $Q\in\mathbb{P}(\Omega)$ such that
$Q(l):=\sum_{m\in\Theta}P(l,m)$
for every $l\in\Omega$. ∎
@@ 5 Postulates of quantum mechanics
In this section, we review the central postulates of (the conventional) quantum mechanics. Specifically, we here refer to the postulates of the conventional quantum mechanics in the form presented in Nielsen and Chuang [30, Chapter 2], with slight modification. Note that their original postulates were given as postulates of (the conventional) quantum mechanics for a finite-dimensional quantum system, i.e., a quantum system whose state space is a finite-dimensional complex Hilbert space, since their book is a textbook of the field of quantum computation and quantum information where finite-dimensional quantum systems are typical. In contrast, we do not impose any restrictions on the dimensionality of the underlying state spaces in the central postulates of (the conventional) quantum mechanics of the form presented in what follows.
@@
Associated to any isolated physical system is a separable complex Hilbert space known as the state space of the system. The system is completely described by its state vector, which is a unit vector in the system’s state space. ∎
@@
The evolution of a closed quantum system is described by a unitary transformation. That is, the state $|\Psi_{1}\rangle$ of the system at time $t_{1}$ is related to the state $|\Psi_{2}\rangle$ of the system at time $t_{2}$ by a unitary operator $U$ on the state space, which depends only on the times $t_{1}$ and $t_{2}$, in such a way that $|\Psi_{2}\rangle=U|\Psi_{1}\rangle$. ∎
@@
A quantum measurement is described by a collection $\{M_{m}\}_{m\in\Omega}$ of measurement operators which is a finite collection of bounded operators on the state space of the system being measured, satisfying the completeness equation
$\sum_{m\in\Omega}M_{m}^{\dagger}M_{m}=I,$
where $M_{m}^{\dagger}$ denotes the adjoint of the bounded operator $M_{m}$, and $I$ denotes the identity operator on the state space.
@@
    The set of possible outcomes of the measurement equals the non-empty finite set $\Omega$. If the state of the quantum system is $|\Psi\rangle$ immediately before the measurement then the probability that result $m$ occurs is given by
@@
In this paper as well as our former works [36, 39, 40, 42, 43, 45, 46, 47, 49, 50], we keep Postulates 1, 2, and 3 in their original forms without any modifications. The principle of typicality, Postulate 5 below, is proposed as a refinement of Postulate 4 to replace it, based on the notion of Martin-Löf randomness with respect to a probability measure.
@@ 6 The principle of typicality: Our world is typical
Tadaki [42] reformulated the original framework of MWI by Everett [22] in a form of mathematical rigor from a modern point of view. The point of this rigorous treatment of the framework of MWI is the use of the notion of probability measure representation and its induction of probability measure, as presented in Section 2.2. Tadaki [42] then introduced the principle of typicality in this rigorous framework. The principle of typicality is an operational refinement of Postulate 4 based on the notion of Martin-Löf randomness with respect to a probability measure given in Definition 2.1. In what follows, we review this rigorous framework of MWI based the principle of typicality. See Tadaki [47] for the details of this rigorous framework of MWI and its validity. Also, for the details of the difference between this framework of MWI and that of the original MWI [22], see Tadaki [47].
Now, according to Tadaki [47], let us introduce the setting of MWI in terms of our terminology in a form of mathematical rigor. Let $\mathcal{S}$ be an arbitrary quantum system with state space $\mathcal{H}$ of an arbitrary dimension. Consider a measurement over $\mathcal{S}$ described by arbitrary measurement operators $\{M_{m}\}_{m\in\Omega}$ satisfying the completeness equation,
@@
Here, $\Omega$ is the set of all possible outcomes of the measurement. [^1] The set $\Omega$ is non-empty and finite, and therefore it is an alphabet. Let $\mathcal{A}$ be an apparatus performing the measurement described by $\{M_{m}\}_{m\in\Omega}$, which is a quantum system with state space $\overline{\mathcal{H}}$. According to Postulates 1, 2, and 3, the measurement process of the measurement described by the measurement operators $\{M_{m}\}_{m\in\Omega}$ is described by a unitary operator $U$ such that
@@
for every $|\Psi\rangle\in\mathcal{H}$. This is due to von Neumann [53, Section VI.3] in the case where the measurement operators $\{M_{m}\}_{m\in\Omega}$ form a PVM in $\mathcal{H}$ such that $M_{m}$ is of rank $1$, i.e., $\dim M_{m}(\mathcal{H})=1$, for every $m\in\Omega$. For the derivation of (5) in the general case, see Tadaki [47, Section 7.1]. Actually, $U$ describes the interaction between the system $\mathcal{S}$ and the apparatus $\mathcal{A}$. The vector $|\Phi_{\mathrm{init}}\rangle\in\overline{\mathcal{H}}$ is the initial state of the apparatus $\mathcal{A}$, and $|\Phi[m]\rangle\in\overline{\mathcal{H}}$ is a final state of the apparatus $\mathcal{A}$ for each $m\in\Omega$, with $\langle\Phi[m]|\Phi[m^{\prime}]\rangle=\delta_{m,m^{\prime}}$. For every $m\in\Omega$, the state $|\Phi[m]\rangle$ indicates that the apparatus $\mathcal{A}$ records the value $m$ as the measurement outcome. By the unitary interaction (5) as a measurement process, a correlation (i.e., entanglement) is generated between the system $\mathcal{S}$ and the apparatus $\mathcal{A}$.
@@
For aiming at deriving Postulate 4, the original MWI [22] assigned “weight” to each of worlds. In our rigorous framework of MWI, we introduce a probability measure on the set of all worlds in the following manner: First, we introduce a probability measure representation on the set of prefixes of worlds, i.e., the set $\Omega^{*}$ in this case. This probability measure representation is given by a function $r\colon\Omega^{*}\to[0,1]$ with
@@
which is the square of the norm of each vector $(M_{m_{1}}|\Psi_{1}\rangle)\otimes\dots\otimes(M_{m_{n}}|\Psi_{n}\rangle)\otimes|\Phi[m_{1}\dots m_{n}]\rangle$ in the superposition (6). Using the completeness equation (4), it is easy to check that $r$ is certainly a probability measure representation over $\Omega$. We call the probability measure representation $r$ the probability measure representation for the prefixes of worlds. We then adopt the probability measure induced by the probability measure representation $r$ for the prefixes of worlds as the probability measure on the set of all worlds.
@@
Consider an arbitrary quantum system $\mathcal{S}$ and a measurement over $\mathcal{S}$ described by arbitrary measurement operators $\{M_{m}\}_{m\in\Omega}$, where the measurement process is described by (5) as an interaction of the system $\mathcal{S}$ with an apparatus $\mathcal{A}$. We suppose the following situation:
@@
    For each $n\in\mathbb{N}^{+}$, the system $\mathcal{S}_{n}$ is prepared in a state $|\Psi_{n}\rangle$, [^2] In Definition 6.1, all $|\Psi_{n}\rangle$ are not required to be an identical state. In the applications of the principle of typicality in this paper, which are presented in later sections, all $|\Psi_{n}\rangle$ are chosen to be an identical state, however. while the apparatus $\mathcal{A}_{n}$ is prepared in a state $|\Phi_{\mathrm{init}}\rangle$, and then the measurement described by $\{M_{m}\}_{m\in\Omega}$ is performed over $\mathcal{S}_{n}$ by interacting it with the apparatus $\mathcal{A}_{n}$ according to the unitary time-evolution (5).
@@
As mentioned above, the original MWI by Everett [22] aimed to derive Postulate 4 (i), the Born rule, from the remaining postulates. However, it would seem impossible to do this for several reasons (see Tadaki [47] for these reasons). Instead, it is appropriate to introduce an additional postulate in our rigorous framework of MWI developed above, in order to overcome the defect of the original MWI and to make quantum mechanics operationally perfect. Thus, we put forward the principle of typicality as follows.
@@
We assume the notation and notions introduced and defined in Definition 6.1, and consider the same setting of measurements as considered in Definition 6.1. Let $\omega\in\Omega^{\infty}$ be our world. Then $\omega$ is Martin-Löf random with respect to the probability measure $\mu_{p}$ induced by the probability measure representation $p\colon\Omega^{*}\to[0,1]$ for the prefixes of worlds. Moreover, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the composite system $\mathcal{S}_{n}+\mathcal{A}_{n}$, which consists of the $n$th copy $\mathcal{S}_{n}$ of the system $\mathcal{S}$ and the $n$th copy $\mathcal{A}_{n}$ of the apparatus $\mathcal{A}$, immediately after the measurement performed by $\mathcal{A}_{n}$ over $\mathcal{S}_{n}$ is given by
@@
7 Apparatus measured by apparatuses and its confirming point